OpenStudy (anonymous):

Sally shot 42 goals in a netball game, which was 3/5 of the team’s total. How many goals were scored by the team? I thought that 3/5=42 so if I change 3/5 into a % I get 60%. So 60%=42 therefore 10%=7. So 100%= 7x10. Is this correct at all?

OpenStudy (mathstudent55):

The problem is stating that 3/5 of a number is 42. What is the number? Let the number be x, since it's an unknown, and write an equation. (3/5)x = 42 Now solve the equation. Since x is being multiplied by 3/5, multiply both sides by the reciprocal of 3/5, which is 5/3.

OpenStudy (anonymous):

I thought that 3/5=42 so if I change 3/5 into a % I get 60%. So 60%=42 therefore 10%=7. So 100%= 7x10. Is this correct at all?

OpenStudy (butterflydreamer):

@zari3200. Your method is also correct :) Both methods end up with the same answer.

OpenStudy (mathstudent55):

@zari3200 Your idea is correct, but you need to be careful how you write it. 3/5 is a number. 42 is another number. \(\dfrac{3}{5} \ne 42\), but \(\dfrac{3}{5}x = 42\) is the correct equation for this problem. As you stated, 3/5 = 60%, so 60% x = 42 10% x = 7, so 100% x = 70, and x, the number we are looking for, is 70.